OpenStudy (anonymous):

Let m(x)=1/(f(x)^3) What is m'(-1)? f(-1)=-9 f'(-1)=7

OpenStudy (anonymous):

\[m(x)=\frac{1}{f^3(x)}=f^{-3}(x)\]

OpenStudy (anonymous):

\[m'(x)=-3f^{-4}(x)f'(x)\] by the chain rule

OpenStudy (anonymous):

so \[m'(1)=-3\times f^{-4}(1)\times f'(1)\]

OpenStudy (anonymous):

now substitute the numbers to get your answer. it is \[-3(-9)^{-4}\times 7\] or \[\frac{-3\times 7}{(-9)^4}\]

OpenStudy (anonymous):

in the second part why is it -3f^-4(x)f'(x) shouldnt it just be -3f^-4(x)

OpenStudy (anonymous):

no chain rule here. that is why you also need \[f'(1)\]